# TituLLMs: A Family of Bangla LLMs with Comprehensive Benchmarking

### v1

@@ 7 Conclusion
In this study, we present the first pretrained Bangla LLMs, Titulm, trained on $\sim 37b$ tokens by adapting Llama-3.2 models. We extended the tokenizer to incorporate language- and culture-specific knowledge, which also enable faster training and inference. Pretraining data collection remains challenging for languages with low digital representation. To address this, we provide a comprehensive approach, including raw web data collection, translation, and synthetic data generation. Given the lack of LLM-based benchmarking datasets, we developed five datasets comprising $137k$ samples, covering both knowledge and reasoning. The benchmarking dataset includes manually curated samples as well as a novel translation-based (EST) approach. Using these datasets, we benchmarked various LLMs, including Titulm, demonstrating its superior performance in reasoning tasks without instruction tuning. Future work includes collecting larger pretraining datasets and fine-tuning with instruction-based datasets. We have made our models publicly available, and to support reproducibility, we plan to release training recipes and benchmarking datasets.
@@ 8 Limitations
We do not anticipate any ethical concerns in this study. All datasets used were collected from publicly available sources, ensuring compliance with ethical research standards. No personally identifiable information (PII) was gathered or utilized in the development of our models. While we do not foresee any potential risks arising from the outcomes of this study, we strongly encourage users of the released models to adhere to responsible AI usage guidelines. This includes avoiding the generation or dissemination of harmful, misleading, or biased content and ensuring that the models are employed in ethical and socially beneficial applications.
@@ Appendix B Tokenizer Details
Appendix C Expressive Semantic Translation (EST)
C.1 Expressive Semantic Translation
@@ C.2 Evaluation
Evaluation metrics: We employ the BLEU Score Papineni et al. (2002), SacreBLEU Score Post (2018), BERT Score (F1) Zhang et al. (2020), and an LLM-Based Evaluation.
@@
Table 5: Comparison of different translation models based on BLEU, SacreBLEU (SBLEU), BERTScore (BS) F1, and LLM scores.
Appendix D Conversation Data Generation Prompt
@@
Appendix E Benchmarking Datasets
Figure 6: Distribution of an benchmarking dataset totaling $\sim 132k$ entries.
In Figure 6, we present the overall distribution of the benchmarking dataset, which consists of approximately 132k entries.
The dataset is composed of multiple subsets of the benchmarking set, including Bangla MMLU (87,869 entries), Piqa BN (17,177 entries), CommonsenseQA BN (10,962 entries), OpenBookQA BN (5,944 entries), and BoolQ BN (1,976 entries). This distribution highlights the significant dominance of the Bangla MMLU subset within the overall evaluation dataset.
Table 7 provides a detailed breakdown of the Bangla MMLU dataset, which contains 116,503 questions spanning multiple educational categories. These include University Admission (62,906), Higher Secondary (32,512), Job Exams (12,864), Medical Admission (4,779), and Engineering Admission (3,442). The dataset reflects a diverse range of question types relevant to various levels of academic and professional assessments, making it a comprehensive benchmark for evaluating LLMs in Bangla educational contexts.
Figure 7: The Bangla MMLU dataset comprises a total of 116,503 questions distributed across various educational categories: University Admission (62,906), Higher Secondary (32,512), Job Exams (12,864), Medical Admission (4,779), and Engineering Admission (3,442).

### v3

@@ 7 Conclusion
In this study, we present the first pretrained Bangla LLMs, TituLLMs, trained on $\sim 37b$ tokens by adapting Llama-3.2 models. We extended the tokenizer to incorporate language- and culture-specific knowledge, which also enable faster training and inference. Pretraining data collection remains challenging for languages with low digital representation. To address this, we provide a comprehensive approach, including raw web data collection, translation, and synthetic data generation. Given the lack of LLM-based benchmarking datasets, we developed five datasets comprising $137k$ samples, covering both knowledge and reasoning. The benchmarking dataset includes manually curated samples as well as a novel translation-based (EST) approach. Using these datasets, we benchmarked various LLMs, including TituLLMs, demonstrating its superior performance in reasoning tasks without instruction tuning. Future work includes collecting larger pretraining datasets and fine-tuning with instruction-based datasets.
@@ 8 Limitations
We do not anticipate any ethical concerns in this study. All datasets used were collected from publicly available sources, ensuring compliance with ethical research standards. No personally identifiable information (PII) was gathered or utilized in the development of our models. While we do not foresee any potential risks arising from the outcomes of this study, we strongly encourage users of the released models to adhere to responsible AI usage guidelines. This includes avoiding the generation or dissemination of harmful, misleading content and ensuring that the models are employed in ethical and socially beneficial applications.
@@ Appendix B Tokenizer Details
Appendix C Translation Model (For pertaining data) Evaluation
Scalar Quality Metrics (SQM): We use Scalar Quality Metrics (SQM) Lankford et al. (2022) to perform human-level evaluation of machine translation outputs in a structured and interpretable manner. SQM employs a 0 to 6 scale to assess the quality of each translated segment, considering both the semantic accuracy and grammatical correctness within the context of the document. This scale provides annotators with a simplified alternative to the traditional 0–100 WMT scoring method, while still capturing nuanced quality differences. Key levels on the scale (0, 2, 4, 6) represent distinct degrees of translation quality—from complete nonsense (0) to perfect translations (6)—with intermediate scores (1, 3, 5) allowing for more precise evaluation. By using SQM, we ensure consistent and context-aware human judgments across different translation models.
To evaluate translation models, we have collected 100 sentences from the BBC News Dataset Greene and Cunningham (2006). We have divided the samples into five categories: Business, Entertainment, Politics, Tech, and Sport. We have 20 samples for each category. After that, we provided the translations from multiple models to two human evaluators (non-authors), and they scored the translations coming from each model.
Appendix D Expressive Semantic Translation (EST)
D.1 Expressive Semantic Translation
@@ D.2 Evaluation
Evaluation metrics: We employ the BLEU Score Papineni et al. (2002), SacreBLEU Score Post (2018), BERT Score (F1) Sun et al. (2022), and an LLM-Based Evaluation.
@@
Table 6: Comparison of different translation models.
Appendix E Conversation Data Generation Prompt